The electrical resistance R of a wire varies directly as the length L and inversely as the square of the diameter d. Obtain an equation connecting R,L and d if R =27 when L=3 and d=2. Calculate (a) value of R when L=8, d=3
(b) value of d when R=16,L=18

Answer by Alan3354(69443)

The electrical resistance R of a wire varies directly as the length L and inversely as the square of the diameter d.
Obtain an equation connecting R,L and d if R =27 when L=3 and d=2.
----
R = k*L/d^2 ---- k = a constant
27 = k*3/4
k = 36
=============================
Calculate (a) value of R when L=8, d=3
Sub, do arithmetic
=========================
(b) value of d when R=16,L=18
Sub, do arithmetic